OpenStudy (anonymous):

Ok so I'm stuck on this problem and I really need help. For 1 sale, an employee earns 1 token. For 2 sales, an employee earns 8 tokens. For 3 sales, an employee earns 17 tokens. For 4 sales, an employee earns 28 tokens. Which recursive equation represents the pattern?

OpenStudy (anonymous):

I think the answer is either an = an - 1 + (n + 2) or an = an - 1 + 2n

OpenStudy (anonymous):

i hate these cuz they are very confusing.

OpenStudy (anonymous):

yes me too lol

OpenStudy (anonymous):

okay so is this geometric or arithmetic?

OpenStudy (anonymous):

arithmetic it's for my financial algebra class

OpenStudy (anonymous):

okay so if its arithmetic you use the recursive formula for those.

OpenStudy (anonymous):

Igbasallote yes

OpenStudy (anonymous):

yes that is what I'm trying to figure out and I'm stuck with on of those two answers listed above

OpenStudy (anonymous):

*one sorry

OpenStudy (anonymous):

28=17+11 A(n)=A(n-1)+d

OpenStudy (anonymous):

ok that sort of helps but these are my choices an = an - 1 + (n + 2) an = an - 1 + (n + 3) an = an - 1 + 2n an = an - 1 + 2n + 3

OpenStudy (anonymous):

A(n) = A(n-1) + (5+2(n-1))

OpenStudy (inkyvoyd):

7-9-11

OpenStudy (inkyvoyd):

@Bree309 , since you have formulas given, just substitute from the data you have - the one that works for all A(n) is what you ar elooking for.

OpenStudy (anonymous):

thats what i wanted to say.

OpenStudy (anonymous):

Yeah I think I have it Thanks guys :)

OpenStudy (anonymous):

feuwwww